OpenStudy (anonymous):

Calculate the area bound by the curve y=x^3-6x^2+11x-6 and the x-axis.

OpenStudy (anonymous):

find the roots of this equation and integrate it from one root to the other

OpenStudy (anonymous):

\[\int\limits_{2}^{1}x^3-6x^2+11x-6 dx +\int\limits_{2}^{3}-(x^3-6x^2+11x-6)dx\] is this correct?

OpenStudy (anonymous):

this can be written (x-3)(x-2)(x-1) integrate from 1 to 3 directly

OpenStudy (anonymous):

\[\int\limits_{1}^{3}x^3 -6x^2 +11x -6\]

OpenStudy (anonymous):

thats how the graph looks its the same on the paper it cuts the x-axis and 3 and 1

OpenStudy (anonymous):

or \[\int\limits_{1}^{2}x^3-6x^2+11x-6 + \int\limits_{2}^{3}x^3-6x^2+11x-6 \]

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\] yeah thats how its suppose to look i have it on paper so for the 2nd integral why'd you put it without a negative sign?

OpenStudy (anonymous):

u dont need a negative sign

OpenStudy (anonymous):

okay thanks him1618!

OpenStudy (anonymous):

if u put a negative sign itll give you the sum of areas and not the algebraic area

OpenStudy (anonymous):

and all i need to do is solve the integrations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or cancel out the 2's and say \[\int\limits_{1}^{3}x^3-6x^2+11x-6 dx\]

OpenStudy (anonymous):

u can do that as well

OpenStudy (anonymous):

and i would get the same answer? okay thanks so much!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

bt here we knew the nature and location of all the roots so we could do that...if your roots were big messy numbers it wldve been wise to just break the integral up

OpenStudy (anonymous):

okay thanks i know what your talking about thanks for the tip!

OpenStudy (anonymous):

no probsd

OpenStudy (anonymous):

lols what?

OpenStudy (anonymous):

u such a noobs him when they say "area", they mean area

OpenStudy (anonymous):

is this wrong?

OpenStudy (anonymous):

u do need the negative for the 2 to 3 region

OpenStudy (anonymous):

or calculate it normally then take the absolute value of it

OpenStudy (anonymous):

is the area 0 because i got 0

OpenStudy (anonymous):

no, the reason you got zero was because you didnt take the absolute value of the stuff below the x axis, and it cancelled with the with the area above. Nevertheless, how can it be zero :|, obvious from the picture that its not zero